OpenStudy (anonymous):

for mathematics distracted a question: i prove 2=1 ;)

OpenStudy (anonymous):

a=b a²=ab a²-b²=ab-b² (a+b)(a-b)=b(a-b) a+b=b b+b=b 2b=b 2=1 ;) error? where? ;)

OpenStudy (amistre64):

dividing by zero is never a good idea

OpenStudy (amistre64):

but thats for a different one isnt it lol

OpenStudy (amistre64):

a-b is still 0 tho

OpenStudy (anonymous):

error -> (a-b) divise different zero.

OpenStudy (amistre64):

so 0 = 0 is a true statment, the error is in dividing of the a-b yep

OpenStudy (anonymous):

amistre64 alert person ! :)